40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 16 23:15:44 2014, Jenne, Update, LSC, IFO alignment recovery Note_061614_230723_0.jpg
    Reply  Tue Jun 17 17:14:14 2014, Manasa, Update, LSC, IFO alignment recovery 
       Reply  Tue Jun 17 21:49:13 2014, Jenne, Update, LSC, IFO alignment recovery 
    Reply  Wed Jun 18 11:57:44 2014, not Jenne, Update, LSC, IFO alignment recovery 
Message ID: 10047     Entry time: Mon Jun 16 23:15:44 2014     Reply to this: 10051   10055
Author: Jenne 
Type: Update 
Category: LSC 
Subject: IFO alignment recovery 

I noticed today, and Rana said that he saw Saturday, that the MC refl value when the MC is unlocked is unusually high.  It typically goes to about 4.5 V, but now is going up to 6.5V.  Since the PMC output is the same as usual (max seen has been about 0.82 today), something must have happened between the PMC and the IMC. 

Late last week, EricG and Nichin were looking at things on the AS table.  Was anything touched on the AS table?  Was anything touched on the PSL table?  'Fess up please, so that we can pinpoint what the change was.

 

Also, this afternoon, I touched up the MC alignment a bit, although it still needs work (I've asked Manasa to look at it tomorrow).  Rana centered the WFS to my MC alignment (this will need to be redone after the MC is truely aligned), and we turned the WFS on.  I also locked both arms individually, and locked MICH and PRMI sideband.  The PRMI wasn't especially stable unless I turned on the POP ASC.  I assume (hope) that this is just because I was doing it during the day, and not because there is something actually different about the PRMI since the computer meltdown.

 

Rana and I also took some notes on things that need to be done, starting tomorrow (the first line and the yellow line are scribbles):

Note_061614_230723_0.jpg

ELOG V3.1.3-